OpenStudy (anonymous):

What is the sum of the geometric sequence -1, 6, -36, … if there are 7 terms

OpenStudy (anonymous):

\[S7=-1.((-6)^{7}-1)/(-6-1)\]

OpenStudy (anonymous):

{-1, 6, -36, 216, -1296, 7776, -46656} Sum=-39991

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Or you can doit as suggested above as \[ S_7=-1\frac{1-(-6)^7}{1+6}=-3999 \]